OpenStudy (anonymous):

The potential energy of a catapult was completely converted into kinetic energy by releasing a small stone with a mass of 20 grams. The velocity of the stone when it reached the target was 15.0 meters/second. What was the value of the kinetic energy of the stone when it reached the target?

OpenStudy (thuru7):

\[E _{k}=\frac{ 1 }{ 2 }mv ^{2}\] = \[\frac{ 1 }{ 2 }\times \frac{ 20 }{ 1000 }\times15^{2}\] \[0.01\times15^{2} =2.25J\]